OpenStudy (anonymous):

Suppose we have a bag with 8 slips of paper in it. Six of these have a 1 on them and the other two have a 3 on them. What is the expected value of the number shown when we draw a single slip of paper?

OpenStudy (anonymous):

easy way to do it is this pick all 8 you get a total of \(6+2\times 3=12\)

OpenStudy (ineedhelpfast12):

12?

OpenStudy (anonymous):

averaged over the 8 picks, the expected value is \(\frac{12}{8}=\frac{3}{2}\)

OpenStudy (anonymous):

Thanks! That helps a lot!

OpenStudy (anonymous):

yw